# On the mixed UDA states and additivity

### v1

@@ Appendix B Proof of tripartite UDA states
$\sigma=$ $\oplus_{l=1}^{j_{1}-1}p_{l}\rho_{A_{2}A_{3}}\oplus T_{1}\oplus_{l=j_{1}+1}^{j_{2}-1}p_{l}\rho_{A_{2}A_{3}}$
$\oplus T_{2}\oplus_{l=j_{2}+1}^{d_{1}}p_{l}\rho_{A_{2}A_{3}},$
@@
$C_{1}:=\begin{bmatrix}pq_{1}&py&pz&0\\ py^{*}&pq_{2}&px&s_{27}\\ pz^{*}&px^{*}&pq_{3}&0\\ 0&s_{27}^{*}&0&(1-p)q_{3}\end{bmatrix}\geq 0,$ (39)
$C_{2}:=\begin{bmatrix}pq_{1}&py&pz&0\\ py^{*}&pq_{2}&px&0\\ pz^{*}&px^{*}&pq_{3}&s_{36}\\ 0&0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}\geq 0,$ (40)
@@
$\begin{bmatrix}pq_{2}&px&s_{27}\\ px^{*}&pq_{3}&0\\ s_{27}^{*}&0&(1-p)q_{2}\end{bmatrix},\begin{bmatrix}pq_{2}&px&0\\ px^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}$
@@
$\begin{bmatrix}pq_{1}&py&0\\ py^{*}&pq_{2}&s_{27}\\ 0&s_{27}^{*}&(1-p)q_{3}\end{bmatrix},\begin{bmatrix}pq_{3}&0&s_{36}\\ 0&(1-p)q_{1}&(1-p)y\\ s_{36}^{*}&(1-p)y^{*}&(1-p)q_{2}\end{bmatrix}$.
@@
$\begin{bmatrix}pq_{1}&pz&0\\ pz^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix},\begin{bmatrix}pq_{2}&0&s_{27}\\ 0&(1-p)q_{1}&(1-p)z\\ s_{27}^{*}&(1-p)z^{*}&(1-p)q_{3}\end{bmatrix}$.
@@
$D_{1}:=\begin{bmatrix}q_{2}&x\\ x^{*}&q_{3}\end{bmatrix},\quad B_{3}:=\begin{bmatrix}pq_{2}&s_{23}\\ s_{23}^{*}&pq_{3}\end{bmatrix}\geq 0,$ (43)
$B_{4}:=\begin{bmatrix}(1-p)q_{2}&-s_{23}+x\\ (-s_{23}+x)^{*}&(1-p)q_{3}\end{bmatrix}\geq 0.$ (44)
@@
$\begin{bmatrix}pq_{2}&s_{23}&s_{27}\\ s_{23}^{*}&pq_{3}&0\\ s_{27}^{*}&0&(1-p)q_{3}\end{bmatrix},\begin{bmatrix}pq_{2}&s_{23}&0\\ s_{23}^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}$ (47)
@@
$\xi:=$ $p\rho_{A_{2}A_{3}}-BV^{\dagger}VB^{\dagger}$ (50)
$=$ $p\rho_{A_{2}A_{3}}-\mathop{\rm diag}\{0,0,\cdots,0,y\}\geq 0,$
@@
(ii) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=4$, then the state $\sigma$ in (49) is also compatible with $\rho$. Further $\sigma\geq 0$ holds since (50) is derived by ${\cal R}(\mathop{\rm diag}\{0,0,\cdots,0,y\})\subseteq{\cal R}(\rho_{A_{2}A_{3}})$ and Lemma 3.   $\sqcap\sqcup$
@@ Appendix C Proof four-qubit UDA states
$\rho=$ $\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$
$=$ $\begin{bmatrix}0&0&0&0\\ 0&p_{2}&a&0\\ 0&a^{*}&p_{3}&0\\ 0&0&0&0\end{bmatrix}\otimes\begin{bmatrix}q_{1}&y&z&0\\ y^{*}&q_{2}&x&0\\ z^{*}&x^{*}&q_{3}&0\\ 0&0&0&0\end{bmatrix}$.
@@
${\cal R}(\rho_{A_{i}A_{i+1}})=\text{span}\{m_{i}|00\rangle+n_{i}|01\rangle,\;g_{i}|00\rangle+h_{i}|10\rangle\},$
@@
(iii.a) ${\cal R}(\rho_{A_{1}A_{2}})$ and ${\cal R}(\rho_{A_{3}A_{4}})$ follow (C). Then $|a_{i}\rangle$’s in (53) are
@@
(iii.c) ${\cal R}(\rho_{A_{1}A_{2}})$ follows (56) and ${\cal R}(\rho_{A_{3}A_{4}})$ follows (C), up to system permutation of $A_{1}A_{2}$ and $A_{3}A_{4}$. We obtain that $S=0$ in TABLE 2, and $\chi=ASA^{\dagger}=0$. Therefore, $\rho$ is 2-UDA. This completes the proof.   $\sqcap\sqcup$
${\cal R}(\rho_{A_{1}A_{2}})$ and ${\cal R}(\rho_{A_{3}A_{4}})$ follow (56). ${\cal R}(\rho_{A_{1}A_{2}})$ follows (56) and ${\cal R}(\rho_{A_{3}A_{4}})$ follows (C)
@@ Appendix D Proof of $n$-qubit UDA states
$\sigma_{A_{1}A_{2}A_{3}}=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}},$
$\sigma_{A_{1}A_{4}A_{5}}=\rho_{A_{1}}\otimes\rho_{A_{4}A_{5}}.$
@@
In order to make $\sigma\geq 0$, we choose nonzero Hermitian $Y$ satisfying ${\cal R}(Y)\subseteq{\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, $Y_{A_{2}A_{3}}=Y_{A_{4}A_{5}}=0$ and $\lVert Y\rVert$ small enough. By the same analysis as Lemma 16, $\sigma$ is congruent to $I_{4}\oplus 0_{12}\oplus(h_{2}\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}-Y^{\dagger}Y)$, which is positive semidefinite. We show the expression of $Y$. Since ${\cal R}(Y)\subseteq{\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, we have $Y=ASA^{\dagger}$, where $S=(s_{jk})$ for $j,k=1,2,3,4$ is Hermitian, and $A=(|a_{1}\rangle,|a_{2}\rangle,|a_{3}\rangle,|a_{4}\rangle)$. Here $|a_{i}\rangle$’s are the base of the vector space ${\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$. Depending on the range of ${\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, this is derived by three cases. First, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ both follow (C). Then $|a_{i}\rangle$’s are given in (57)-(60), and $S$ is diagonal with $s_{11}=x$, $s_{22}=-\frac{|m_{3}|^{2}+|n_{3}|^{2}}{|g_{3}|^{2}+|h_{3}|^{2}}x$, $s_{33}=-\frac{|m_{1}|^{2}+|n_{1}|^{2}}{|g_{1}|^{2}+|h_{1}|^{2}}x$, $s_{44}=-\frac{(|m_{1}|^{2}+|n_{1}|^{2})(|m_{3}|^{2}+|n_{3}|^{2})}{(|g_{1}|^{2}+|h_{1}^{2}|)(|g_{3}|^{2}+|h_{3}^{2}|)}x$, for $x>0$ small enough. Second, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ both follow (56). Then $|a_{1}\rangle=z_{1}z_{3}|00\rangle$, $|a_{2}\rangle=z_{1}|00\rangle\otimes(u_{3}|01\rangle+v_{3}|10\rangle)$, $|a_{3}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes z_{3}|00\rangle$, $|a_{4}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(u_{3}|01\rangle+v_{3}|10\rangle)$. We choose $s_{12}=(|u_{1}|^{2}+|v_{1}|^{2})x$, $s_{34}=-|z_{1}|^{2}x$, and other elements of $S$ are zero, for $x>0$ small enough. Third, up to system permutation of $A_{2}A_{3}$ and $A_{4}A_{5}$, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ follow (56) and (C), respectively. Then $|a_{1}\rangle=z_{1}|00\rangle\otimes(m_{3}|00\rangle+n_{3}|01\rangle)$, $|a_{2}\rangle=z_{1}|00\rangle\otimes(g_{3}|00\rangle+h_{3}|10\rangle)$, $|a_{3}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(m_{3}|00\rangle+n_{3}|01\rangle)$, and $|a_{4}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(g_{3}|00\rangle+h_{3}|10\rangle)$. We choose $s_{13}=-(|g_{3}|^{2}+|h_{3}|^{2})x$, $s_{24}=(|m_{3}|^{2}+|n_{3}|^{2})x$ and other elements of $S$ are zero, for $x>0$ small enough. To conclude, we can always find a nonzero $Y$ in (67) such that $\sigma$ is 2-compatible with $\rho$. This completes the proof.   $\sqcap\sqcup$

### v2

@@ Appendix B Proof of tripartite UDA states
$\sigma=$ $\oplus_{l=1}^{j_{1}-1}p_{l}\rho_{A_{2}A_{3}}\oplus T_{1}\oplus_{l=j_{1}+1}^{j_{2}-1}p_{l}\rho_{A_{2}A_{3}}\oplus T_{2}\oplus_{l=j_{2}+1}^{d_{1}}p_{l}\rho_{A_{2}A_{3}},$
@@
$C_{1}:=\begin{bmatrix}pq_{1}&py&pz&0\\ py^{*}&pq_{2}&px&s_{27}\\ pz^{*}&px^{*}&pq_{3}&0\\ 0&s_{27}^{*}&0&(1-p)q_{3}\end{bmatrix}\geq 0,\quad C_{2}:=\begin{bmatrix}pq_{1}&py&pz&0\\ py^{*}&pq_{2}&px&0\\ pz^{*}&px^{*}&pq_{3}&s_{36}\\ 0&0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}\geq 0,$ (35)
@@
$\begin{bmatrix}pq_{2}&px&s_{27}\\ px^{*}&pq_{3}&0\\ s_{27}^{*}&0&(1-p)q_{2}\end{bmatrix},\quad\begin{bmatrix}pq_{2}&px&0\\ px^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}$
@@
$\begin{bmatrix}pq_{1}&py&0\\ py^{*}&pq_{2}&s_{27}\\ 0&s_{27}^{*}&(1-p)q_{3}\end{bmatrix},\quad\begin{bmatrix}pq_{3}&0&s_{36}\\ 0&(1-p)q_{1}&(1-p)y\\ s_{36}^{*}&(1-p)y^{*}&(1-p)q_{2}\end{bmatrix}$.
@@
$\begin{bmatrix}pq_{1}&pz&0\\ pz^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix},\quad\begin{bmatrix}pq_{2}&0&s_{27}\\ 0&(1-p)q_{1}&(1-p)z\\ s_{27}^{*}&(1-p)z^{*}&(1-p)q_{3}\end{bmatrix}$.
@@
$D_{1}:=\begin{bmatrix}q_{2}&x\\ x^{*}&q_{3}\end{bmatrix},\quad B_{3}:=\begin{bmatrix}pq_{2}&s_{23}\\ s_{23}^{*}&pq_{3}\end{bmatrix}\geq 0,\quad B_{4}:=\begin{bmatrix}(1-p)q_{2}&-s_{23}+x\\ (-s_{23}+x)^{*}&(1-p)q_{3}\end{bmatrix}\geq 0.$ (38)
@@
$\begin{bmatrix}pq_{2}&s_{23}&s_{27}\\ s_{23}^{*}&pq_{3}&0\\ s_{27}^{*}&0&(1-p)q_{3}\end{bmatrix},\quad\begin{bmatrix}pq_{2}&s_{23}&0\\ s_{23}^{*}&pq_{3}&s_{36}\\ 0&s_{36}^{*}&(1-p)q_{2}\end{bmatrix}$ (41)
@@
$\xi:=p\rho_{A_{2}A_{3}}-BV^{\dagger}VB^{\dagger}=p\rho_{A_{2}A_{3}}-\mathop{\rm diag}\{0,0,\cdots,0,y\}\geq 0,$
@@
(ii) If $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=4$, then the state $\sigma$ in (43) is also compatible with $\rho$. Further $\sigma\geq 0$ holds since (B) is derived by ${\cal R}(\mathop{\rm diag}\{0,0,\cdots,0,y\})\subseteq{\cal R}(\rho_{A_{2}A_{3}})$ and Lemma 3.   $\sqcap\sqcup$
The proof of Lemma 16 is shown below.
Proof.
The first fact is proved by showing the range of $\rho$ has a pure state $|\gamma\rangle$ LU equivalent to the quasi-GHZ state $a|000\rangle+b|111\rangle$ with $ab\neq 0$. More details will be given in the proof of Proposition 26. Suppose $\mathop{\rm rank}(\rho)=6$ and $\mathop{\rm rank}(\rho_{A_{2}A_{3}})=3$. One can show that ${\cal R}(\rho)={\cal R}(\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}})$ with full-rank $\rho_{A_{1}}$. Let $|\psi\rangle\in\ker({\cal R}(\rho_{A_{2}A_{3}}))$ and $\Pi=I_{A_{1}}\otimes|\psi\rangle\!\langle\psi|$. Then $0=\langle\psi|\rho_{A_{2}A_{3}}|\psi\rangle=\mathop{\rm Tr}[(I_{A_{1}}\otimes|\psi\rangle\!\langle\psi|)\rho]=\mathop{\rm Tr}(\Pi\rho)$. Since $\Pi\rho\Pi\geq 0$ and $\mathop{\rm Tr}(\Pi\rho\Pi)=\mathop{\rm Tr}(\Pi\rho)=0$, we have $\Pi\rho\Pi=0$. This implies that ${\cal R}(\rho)=\mathop{\rm supp}(\rho)\subseteq{\cal H}_{A_{1}}\otimes{\cal R}(\rho_{A_{2}A_{3}})={\cal R}(\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}})$. From $\mathop{\rm Dim}({\cal R}(\rho))=\mathop{\rm Dim}({\cal R}(\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}))$, we have ${\cal R}(\rho)={\cal R}(\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}})$. Since $\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}}$ is not 2-UDA, we have $\rho$ is not 2-UDA. This completes the proof.   $\sqcap\sqcup$
@@ Appendix C Proof four-qubit UDA states
$\rho=\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}=\begin{bmatrix}0&0&0&0\\ 0&p_{2}&a&0\\ 0&a^{*}&p_{3}&0\\ 0&0&0&0\end{bmatrix}\otimes\begin{bmatrix}q_{1}&y&z&0\\ y^{*}&q_{2}&x&0\\ z^{*}&x^{*}&q_{3}&0\\ 0&0&0&0\end{bmatrix}$.
@@
${\cal R}(\rho_{A_{i}A_{i+1}})=\text{span}\{m_{i}|00\rangle+n_{i}|01\rangle,\;g_{i}|00\rangle+h_{i}|10\rangle\},$ (48)
@@
(iii.a) ${\cal R}(\rho_{A_{1}A_{2}})$ and ${\cal R}(\rho_{A_{3}A_{4}})$ follow (48). Then $|a_{i}\rangle$’s in (46) are
@@
(iii.c) ${\cal R}(\rho_{A_{1}A_{2}})$ follows (49) and ${\cal R}(\rho_{A_{3}A_{4}})$ follows (48), up to system permutation of $A_{1}A_{2}$ and $A_{3}A_{4}$. We obtain that $S=0$ in TABLE 2, and $\chi=ASA^{\dagger}=0$. Therefore, $\rho$ is 2-UDA. This completes the proof.   $\sqcap\sqcup$
${\cal R}(\rho_{A_{1}A_{2}})$ and ${\cal R}(\rho_{A_{3}A_{4}})$ follow (49). ${\cal R}(\rho_{A_{1}A_{2}})$ follows (49) and ${\cal R}(\rho_{A_{3}A_{4}})$ follows (48)
@@ Appendix D Proof of $n$-qubit UDA states
$\sigma_{A_{1}A_{2}A_{3}}=\rho_{A_{1}}\otimes\rho_{A_{2}A_{3}},$ (57)
$\sigma_{A_{1}A_{4}A_{5}}=\rho_{A_{1}}\otimes\rho_{A_{4}A_{5}}.$ (58)
@@
In order to make $\sigma\geq 0$, we choose nonzero Hermitian $Y$ satisfying ${\cal R}(Y)\subseteq{\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, $Y_{A_{2}A_{3}}=Y_{A_{4}A_{5}}=0$ and $\lVert Y\rVert$ small enough. By the same analysis as Lemma 18, $\sigma$ is congruent to $I_{4}\oplus 0_{12}\oplus(h_{2}\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}}-Y^{\dagger}Y)$, which is positive semidefinite. We show the expression of $Y$. Since ${\cal R}(Y)\subseteq{\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, we have $Y=ASA^{\dagger}$, where $S=(s_{jk})$ for $j,k=1,2,3,4$ is Hermitian, and $A=(|a_{1}\rangle,|a_{2}\rangle,|a_{3}\rangle,|a_{4}\rangle)$. Here $|a_{i}\rangle$’s are the base of the vector space ${\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$. Depending on the range of ${\cal R}(\rho_{A_{2}A_{3}}\otimes\rho_{A_{4}A_{5}})$, this is derived by three cases. First, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ both follow (48). Then $|a_{i}\rangle$’s are given in (50)-(53), and $S$ is diagonal with $s_{11}=x$, $s_{22}=-\frac{|m_{3}|^{2}+|n_{3}|^{2}}{|g_{3}|^{2}+|h_{3}|^{2}}x$, $s_{33}=-\frac{|m_{1}|^{2}+|n_{1}|^{2}}{|g_{1}|^{2}+|h_{1}|^{2}}x$, $s_{44}=-\frac{(|m_{1}|^{2}+|n_{1}|^{2})(|m_{3}|^{2}+|n_{3}|^{2})}{(|g_{1}|^{2}+|h_{1}^{2}|)(|g_{3}|^{2}+|h_{3}^{2}|)}x$, for $x>0$ small enough. Second, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ both follow (49). Then $|a_{1}\rangle=z_{1}z_{3}|00\rangle$, $|a_{2}\rangle=z_{1}|00\rangle\otimes(u_{3}|01\rangle+v_{3}|10\rangle)$, $|a_{3}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes z_{3}|00\rangle$, $|a_{4}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(u_{3}|01\rangle+v_{3}|10\rangle)$. We choose $s_{12}=(|u_{1}|^{2}+|v_{1}|^{2})x$, $s_{34}=-|z_{1}|^{2}x$, and other elements of $S$ are zero, for $x>0$ small enough. Third, up to system permutation of $A_{2}A_{3}$ and $A_{4}A_{5}$, ${\cal R}(\rho_{A_{2}A_{3}})$ and ${\cal R}(\rho_{A_{4}A_{5}})$ follow (49) and (48), respectively. Then $|a_{1}\rangle=z_{1}|00\rangle\otimes(m_{3}|00\rangle+n_{3}|01\rangle)$, $|a_{2}\rangle=z_{1}|00\rangle\otimes(g_{3}|00\rangle+h_{3}|10\rangle)$, $|a_{3}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(m_{3}|00\rangle+n_{3}|01\rangle)$, and $|a_{4}\rangle=(u_{1}|01\rangle+v_{1}|10\rangle)\otimes(g_{3}|00\rangle+h_{3}|10\rangle)$. We choose $s_{13}=-(|g_{3}|^{2}+|h_{3}|^{2})x$, $s_{24}=(|m_{3}|^{2}+|n_{3}|^{2})x$ and other elements of $S$ are zero, for $x>0$ small enough. To conclude, we can always find a nonzero $Y$ in (60) such that $\sigma$ is 2-compatible with $\rho$. This completes the proof.   $\sqcap\sqcup$